OpenStudy (anonymous):

How would you get the total mass by adding 0.500 g of Ca to 50.0 mL of 1.0 M of HCL?

OpenStudy (rushwr):

Total mass = mass of calcium + mass of HCl We know the mass of Ca so we don't have to find it. We only have to find the mass of HCl We are given with the volume and the concentration of HCl, so use n=CV equation to find the no. of moles. (C= concentration and V= volume , n= no. of moles) We know that no. of moles is equal to mass divided by the molar mass. \[n = \frac{ m }{ M }\] m= mass of HCl M= molar mass Molar mass of HCl = 1+ 35.5 - 36.5gmol^-1 Since we know the molar mass and the no. of moles of HCl, we can find the mass of HCl

OpenStudy (rushwr):

Did u get it ? @radiskull

OpenStudy (anonymous):

Lookin at it right now, wait one.

OpenStudy (rushwr):

And remember here the volume is given in milliliters, so we have to convert it to cubic decimeter to apply it to n= CV equation !

OpenStudy (anonymous):

OK thanks, I'll give it a try.

OpenStudy (rushwr):

Anytym :) if u can give me the answer now I can check it for you !

OpenStudy (anonymous):

Nope I couldn't figure it out, never used this equation before n = CV

OpenStudy (rushwr):

Really? At Least something like \[C = \frac{ n }{ V }\]

OpenStudy (rushwr):

Didn't u still get it ?

OpenStudy (anonymous):

I'll figure it out, thankx. Appreciate it!

OpenStudy (rushwr):

Medal me if u don't mind :) And WELCOME TO OPEN STUDY !

OpenStudy (anonymous):

Thanks Rushwr very kind of you.